OpenStudy (anonymous):

Find the point of intersection of the pair using the method of elimination x-y = 3 2x + y = 3

OpenStudy (mertsj):

If you just add the two equations, the y will be eliminated. Can you do that?

OpenStudy (anonymous):

I don't understand what the question is exactly asking

OpenStudy (katrinakaif):

You have to solve for x and y, concluding with an ordered pair of (x,y)

OpenStudy (karatechopper):

well elimination is where you eliminate one of the variables to be able to solve the equation

OpenStudy (mertsj):

Ok Katrina and karate will help you.

OpenStudy (anonymous):

Can anybody solve it for me? I don't really get what you're trying to say....

OpenStudy (katrinakaif):

Ok. One min. My explanations will be followed.

myininaya (myininaya):

add the equations together (x+2x)+(-y+y)=(3+3) that should help

OpenStudy (katrinakaif):

Your equations are: x - y = 3 2x + y = 3 In both equations "-y" from the first equation, and "+y" from the second equation cancel each other out. Understood?

OpenStudy (karatechopper):

therefore lleaving the equation to be x=3 2x=3

OpenStudy (katrinakaif):

Since they cancel each other, the equation becomes: 3x = 6

OpenStudy (katrinakaif):

Now lets solve for x. Divide both sides for 3, which concludes us with the answer x = 2 Since we have solved for x, our ordered pair would be (2,y)

OpenStudy (karatechopper):

now you take that 2 and plug it into one of the equations

OpenStudy (karatechopper):

im going to take the x-y=3 equation so i plug 2 in for x 2-y=3 subtract 2 on both sides -y=1 divide by -1 on both sides y=1

OpenStudy (katrinakaif):

Now, use what you have found as being x= 2, and plug it into your original equation. Lets plug "2" in the position of "x" in equation one: x - y = 3 (2) - y = 3 Bring the 2 to the "other side" but substracting it from 3 -y = 3 - 2 -y = 1 Divide by -1 and your answer is y = -1

OpenStudy (katrinakaif):

Now since you have solved x and y, your coordinates of intersection is (2, - 1)

OpenStudy (karatechopper):

woops i messed up make that be -1

OpenStudy (karatechopper):

nice work kat!;) medal timne!!!

OpenStudy (katrinakaif):

That will be your final answer. Any questions?

OpenStudy (anonymous):

Thank you so much. God bless <3

hero (hero):

KC and KK helping people with math? Has the world turned upside down?

OpenStudy (katrinakaif):

Anytime Waleed =)

OpenStudy (karatechopper):

@hero what wrong with that? i have always helped people with the math tht i know..

hero (hero):

It's when both of you do it together....that's what makes it strange.

OpenStudy (karatechopper):

???

OpenStudy (karatechopper):

its called TEAMWORK!

hero (hero):

i'm a one man team.

OpenStudy (karatechopper):

ok kool;)

OpenStudy (anonymous):

the more the merrier :D

OpenStudy (karatechopper):

:D

hero (hero):

will only be happier